OpenStudy (needhelpp101):

Translate the problem situation to a system of equations. Do not solve. The difference between two numbers is 1. Three times the smaller number plus the larger number is 29. What are the numbers? A. x - y = 1; x + 3z = 29 B. x - y = 1; 3x + y = 29 C. x - y = 1; x + 3y = 29 D. x + y = 1; x + 3y = 29

OpenStudy (haiitskim):

Umm I think C..

OpenStudy (haiitskim):

Wait no..

OpenStudy (haiitskim):

B

OpenStudy (needhelpp101):

lol ok are you sure?

OpenStudy (anonymous):

answer is c

OpenStudy (anonymous):

since the difference is 1 which is positive...this means x.>y x+3y=29,x-y =1

OpenStudy (anonymous):

@haiitskim....don't give wrong answer

OpenStudy (needhelpp101):

lol thx Manish :)

OpenStudy (anonymous):

alwaz.....welcum

OpenStudy (haiitskim):

@manishsaywali I said c at first.. So technically I didn't :)

OpenStudy (anonymous):

See u hav given that differnce of two no. is 1..which is actually positive nw wen u r taking x-y=1 this mean that x>y that x is greater no. and y is smaller fr the nxt condition u hav given that three times the smaller number plus the larger number is 29. that 3y+x= 29